OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale?

OpenStudy (ivettef365):

the coat price is $100, discount is 15%, first find out what is the price after discount ?

OpenStudy (anonymous):

idk

OpenStudy (ivettef365):

or just plug in the formula you have C = 0.765x + 0.06(0.765x) where x = 100 (price of coat) C = 0.765(100) + 0.06(0.765(100)) C = 76.50 + 0.06(76.50) C = 76.50 + .46 C = 76.96

OpenStudy (ivettef365):

wait I did something wrong C = 76.50 + 0.06(76.50) C = 76.50 + 4.59 C = 81.09

OpenStudy (ivettef365):

now is correct, so you understand what I did ?

OpenStudy (anonymous):

no

OpenStudy (ivettef365):

I just plugged in 100 for the value of x on the formula that was given to you and solved it

OpenStudy (anonymous):

oh ok thanks for ur help

OpenStudy (ivettef365):

yw :)